Fermi National Laboratory
 

Accelerator Update for November 5, 2007 - November 7, 2007

 

Monday November 5

The day shift began with Operations monitoring a Recycler stash (149.8E10), with the Antiproton Source (Pbar) stacking (22.5E10), with the Main Injector (MI), Recycler, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store, and with beam to the SciBooNE / MiniBooNE experiments.

Power Supply experts reported at 9:25 AM that they had resolved the Tevatron (TeV) sector B1 abort problem.

TeV experts began a proton only study at 9:36 AM.

A MI expert began a slipstacking study at 10:57 AM.

At 11:07 AM, the 750 KeV area buncher tripped off. It reset.

At 11:37 AM, an Accumulator RF station (ARF1) tripped off on a transformer over temperature indication and would not reset. An expert investigated and had the station repaired by 11:50 AM.

Operations prepared Linac for an access at 12:58 PM to inspect the LRF3 penetration LCW leak.

Recycler experts accessed the Pelletron at 1:42 PM.

A Linac expert reported 2:05 PM that the LRF3 LCW leak was found under the drift tube cavity's stem box cover on tubing that went to a quadrupole magnet located in the cavity.


Operations began turning off the TeV at 3:17 PM for an access to inspect separators.

Operations reported at 4:01 PM that the Recycler had stashed antiprotons once during the day shift.

Pbar experts began an Accumulator dispersion correction study at 4:12 PM.

At 5:10 PM, a TeV expert reported that he found a bad surge resistor on the D48H separator, but that they wouldn’t have time to fix it nor would they have time to look at the C17V separator. (They will repair these separators during the shutdown on Tuesday.)

Operations reported that the LFR3 LCW leak was repaired by 5:16 PM.

The TeV access ended at 5:43 PM.

Linac resumed sending beam at 6:40 PM.

Pbar experts began a reverse proton study at 8:31 PM.

Operations put the TeV through a dry squeeze at 9:08 PM.

At 9:08 PM, BRF15 tripped off and would not reset. Operators bypassed the station.

Operations put the TeV through a wet squeeze at 9:58 PM.

Operations put the TeV into shot setup at 10:25 PM.

Tuesday November 6

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (140.6E10), with Pbar stacking (0.1E10), and with Sci /MiniBooNE taking beam.

At 12:43 AM, the TeV beam aborted. There was no quench. The abort was caused by a bad beam position in the CDF silicon detector.

Recycler experts began a study at 12:50 AM.

CDF experimenters reset their SVX abort at 2:08 AM. TeV experts began a proton only study.

Operations began preparing the TeV for an access at 3:46 AM.

Cryo system experts began warning up the cryo systems at TeV sectors A4 and B1.

The Recycler studies ended at 5:11 AM.

Operations began turning off Booster and MI at 5:33 AM.

Booster experts accessed their tunnel at 6:51 AM. CDF experimenters accessed their collision hall at the same time.

Pbar experts accessed their tunnels at 7:04 AM.

The Cryo Coordinator reported at 7:50 AM that the TeV sectors A4 and B1 cryo systems had been warmed up to 20 Kelvin.

Operations prepared Linac for access at 9:45 AM.

Operations reported at 10:18 AM that operators had rebooted the MI mid and low level RF systems and the damper system.

At 10:34 AM, a Linac expert reported finding a small LCW leak coming from a LRF3 compression fitting. Water technicians were preparing to fix the leak.

Operations prepared the Meson M01-5 enclosures for access at 10:43 AM.

Vacuum techs reported finding a SwitchYard (SY) vacuum leak on a blower at 101. They isolated the blower and said they would replace it later.

Linac came out of access at noon.

At 12:10 PM, the 750 KeV area buncher tripped off. It reset.

The Pbar work ended and Operations had the machine ready for beam by 12:55 PM.

Vacuum experts reported at 2:15 PM that they had replaced a bad TeV sector F47 vacuum pump O-ring.

MI came out of access and the accelerator was being turned on by 2:33 PM.

The TeV F-Sector and the Transfer Hall came out off access at 3:09 PM.

Booster came out of access at 4:40 PM.

At 4:46 PM, a Pbar Debuncher power supply (D:SEXFV) tripped off on a “Phase Unbalance” indication and would not reset. Operations contacted an expert.

Operations established beam to Sci/MiniBooNE at 5:06 PM.

Recycler experts began a reverse proton study at 5:19 PM.

Operations put Pbar into stacking mode at 6:07 PM.

A controls expert reported at 6:11 PM that he had found and replaced a bad ramp card in Booster.

Pbar experts began a study period at 7:55 PM.

At 9:14 PM, the 750 KeV area buncher tripped off. It reset.

At 9:28 PM, BRF18 tripped off and caused BRF17 and the Booster West Anode Power Supply (WAPS) to trip off. The power supply and BRF17 reset, but BRF18 would not. Operators bypassed the station and brought BRF19 on in its place. Operations called in an RF expert.

Wednesday November 7

The midnight shift began with TeV still in access, with Pbar and Recycler experts conducting studies, and with Sci /MiniBooNE taking beam.

Pbar resumed stacking at 12:54 AM.

At 1:11 AM, a Pbar Accumulator RF station (ARF1) tripped off. It took a while, but operators managed to turn the station back on.

At 2:28 AM, ARF1 tripped off again. Operators investigated and found a small LCW leak. Operations called in an expert.

At 3:09 AM, the 750 KeV area buncher tripped off. It reset.

At 3:18 AM, a Pbar Debuncher power supply (D:H926) tripped off and would not reset. Operations called in an expert.

Pbar resumed stacking at 4:14 AM.

An RF expert reported at 4:17 AM that he replaced a bad fan for ARF1, which brought the station back on. He said the LCW leak was a separate issue.

Operations began preparing the Recycler to stash antiprotons at 7:47 AM.

The Plans for Wednesday

The plans for today are for experts to continue with machine studies while the TeV magnet moves are completed. The moves and other associated work should be completed by this evening. The Run Coordinator wants the instrumentation experts to work on the DCCT problems.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/07/2007 -->    email Fermilab